### v2

# Detection of the papermilling behavior

Abstract
Based on the analysis of the data obtainable from the Web of Science publication and citation database, typical signs of possible papermilling behavior are described, quantified, and illustrated by examples. A MATLAB function is provided for the analysis of the outputs from the Web of Science. A new quantitative indicator – integrity index, or I-index – is proposed for using it along with standard bibliographic and scientometric indicators.
1 Introduction: the good intentions
The original idea of publications and citations indexes was to help scientists in their research work. Eugene Garfield, the creator of the Science Citation Index, wrote:
“In this paper, I propose a bibliographic system for science literature that can eliminate the uncritical citation of fraudulent, incomplete, or obsolete data by making it possible for the conscientious scholar to be aware of criticisms of earlier papers.” [1].
However, there is a proverb that says good intentions pave the road to hell, and the field of mathematics is currently an identified victim of massive scientific misconduct related to manipulations with publications and citations counts considered as indicators of the quality of individual researchers. Clearly, other fields of science suffer similar problems, although perhaps less visible.
2 The current situation: “code red”
In the description of the methodology for preparing the Highly Cited Researchers™  2023 lists, Clarivate™ explains the reasons for excluding the field of mathematics from such list [2, see section “Exceptions and exclusions”]:
“We have chosen to exclude the Mathematics category from our analysis for this year.
The field of Mathematics differs from other categories in ESI. It is a highly fractionated research domain, with few individuals working on a number of specialty topics. The average rate of publication and citation in Mathematics is relatively low, so small increases in publication and citation tend to distort the representation and analysis of the overall field. Because of this, the field of Mathematics is more vulnerable to strategies to optimize status and rewards through publication and citation manipulation, especially through targeted citation of very recently published papers which can more easily become highly cited (top 1% by citation). This not only misrepresents influential papers and people; it also obscures the influential publications and researchers that would have qualified for recognition. The responsible approach now is to afford this category additional analysis and judgement to identify individuals with significant and broad influence in the field.”
It should be added that the lists of highly cited researchers in mathematics contained 99 names in 2014, also 99 names in 2015, 106 names in 2016, 96 names in 2017, 90 names in 2018, 89 names in 2019, 70 names in 2020, 74 names in 2021, 52 names in 2022, and no names after that [3]. These lists are based on the data from the preceding years – for example, the 2022 list is based on the data from 2011 to 2021. In all the lists there are 243 unique names.
It is worth mentioning here, that the fact that publications and citations counts in mathematics are the lowest compared to other sciences is known, and has been used for comparing citation impact in different fields of science [4] and for creating the first multidisciplinary list of highly cited researchers based on their citation counts normalized with respect to mathematics [5].
At the time of writing of the paper [5], the Highly Cited Researchers list (at the time known as Thomson ISI) was compiled based on the citations counts over the duration of the entire career. There was also a category called “Research front”, where were included papers that collected some threshold number of citations within short time (a couple of years only); indeed, this could indicate that such papers might be bringing really new or significant results.
Currently, Clarivate compiles its Highly Cited Researchers (HCR) lists using something similar to those previous “Research front” papers [2, see section “Preliminary candidate selection”]:
“For the Highly Cited Researchers 2023 analysis, the papers surveyed were those published during 2012 to 2022 and which then ranked in the top 1% by citations for their ESI field and year (the definition of a highly cited paper).”
The abbreviation ESI means Essential Science Indicators™  – another Clarivate product.
Instead of the accumulated whole-career citation impact, the “beans” called “highly cited papers” are counted for a relatively short time window. Due to this change of rules, it became possible – exactly as Clarivate noticed – to arrange relatively small numbers of citations for a set of several papers and, as a result, get a particular person to the HCR list. The field of mathematics appeared particularly vulnerable to this type of action, because of its low citations counts per paper.
However, mathematics has the intrinsic property of a usually long delay between the publication of a result and its wide recognition and use, reflected in the number of citations. To artificially increase the number of citations for the produced/selected “beans” within a short time frame, the manipulation of citations is used used, which is done by producing more insignificant papers with the sole purpose of increasing the citation counts for the “beans”.
Apparently, the people who become the “recipients” of a papermill in order to increase their publication counts, also become the “providers” of citations for other recipients through their subsequent papermill publications. This leads to the snowball effect, and is likely the reason for the rapid monotonous growth in publication counts of individuals involved in the papermilling process.
3 The data source
The data for further processing are obtained by exporting necessary search results from the Web of Science (WoS). The necessary steps are as follows.
  • Go to the Web of Science and open Search.
  • Switch to the RESEARCHERS tab, and start typing the last name; in the appearing dropdown list select the necessary last name. Then, similarly, enter the first name (start typing and select from the dropdown list). Eventually, it is possible to add name variants. After entering the surname and the name, click on the “Search” button.
  • A researcher may have one or more affiliations.
    • (a)
      If a researcher has one affiliation, click on the name of the researcher.
    • (b)
      If a researcher has/had more than one affiliation, select proper affiliations using the checkboxes, and click on the “View as combined record” button.
  • In the right pane of the next screen locate and click the “View citation report” button.
  • On the next screen called “Citation Report”, which provides an overview of the researcher’s indicators (publications, citations, H-index, and some other details), click on the link “Export Full Report”, and save all outputs (use the “Records from:” option) using the “Excel File” format.
The above procedure allows exporting up to 1000 records, which sounds like a reasonable number. However, some researchers have more than 1000 publications indexed in the WoS, and in such cases it is necessary to use the “Records from:” option several times with subsequent manual merging of the data from the obtained spreadsheet files.
4 Visual signs of papermilling behavior
On the Web of Science “Citation Report” page there is also the plot titled “Times Cited and Publications Over Time”. The vertical axis on the left is for publications counts, and the vertical axis on the right is for citations counts; these two scales are different.
Figure 1: Typical performance of a conscientious mathematician.
Figure 2: Typical visual signs of possible papermilling in mathematics.
This plot allows for hovering over or clicking on the vertical bars (or on the line) in order to see the number of publications and citations in a particular year.
Two sample outputs are shown in Fig. 1 (researcher R1) and Fig. 2 (researcher R2), and in both cases the researchers are mathematicians who have been listed in the Highly Cited Researchers lists in the past; their career lengths are similar.
Researcher R1 had 196 papers indexed in the Web of Science, with at most 19 items in one year. The publication activity of R1 was slightly increasing before 2011, reaching 18 and 19 publications in 2005 and 2006, after which it started declining. The number of citations of R1 had been growing with a significant delay, reaching its peaks in 2015. The number of R1’s papers before 2011 was 159, with 128 co-authors from 15 countries. In the period of 2011-2023, the number of publications of R1 was 37, with 36 co-authors from 6 countries.
Researcher R2 had 384 papers indexed in the Web of Science, with as many as 45 paper in one year. The publication activity of R2 before 2011 was low, with about 4 papers in one year, 81 papers in total, and with little citation impact; the number of co-authors of R2 before 2011 was 17, from 12 countries.
However, from 2011 on, the number of publications of R2 started rapidly growing, giving the total of 303 publications in 2011–2023, reaching the peak of 45 papers in 2022, and the number of citations started to grow synchronously with the number of publications. The total number of co-authors of R2 during 2011–2023 was 148, from 35 countries.
The described differences between researchers R1 and R2 are obvious. It should be mentioned that the publication pattern of R1 also reflects the generally-known observation that significant results are achieved in the first half of a mathematician’s career, and that the publication activity normally decreases with age. In the case of the researcher R2, one can observe the aforementioned snowball effect typical for the papermilling behavior.
Obviously, in the case of researcher R2, authoring 45 papers in mathematics in one year is not typical, and the synchronous growth of citation counts alongside the growth in number of papers is also atypical. On the contrary, in the case of researcher R1, the continuous growth of citations counts despite the decrease in the number of papers is normal and shows the long-lasting impact of R1’s publications in the first half of the considered time period.
The following features in the plots for publications and citations counts are possible indicators of the papermilling behavior:
  • for a significantly long time period of career, the number of publications and the number of citations of a researcher are low;
  • starting from some particular year, the number of publications is starts rapidly growing (usually monotonically), and reaches a number of papers per year which is unrealistic for a conscientious and responsible scientist;
  • the number of citations exhibits the same rapid growth behavior starting practically at the same time;
  • there is no delay between the rapidly growing number of publications and the rapidly growing number of citations – they increase simultaneously;
  • the number of co-authors (and their affiliations and countries) notably increases when the number of publications starts increasing.
Besides these symptomatic signs of the overall publishing behavior, there are also symptoms that symptoms that require scrutiny of each particular publication.
First, papermilling articles do not bring new ideas, their contribution is minimal (if any), and they are often simply incorrect.
Second, they usually include longer lists of references (with many unnecessary references) and often use bulk citations in the text, like “see [17–22, 34–38]” , or “this was studied in [2]–[15]”. Alternatively, many individual sentences in the introduction contain just one or two references, so the bulk citations are spread across some usually unnecessary text.
5 Basic quantitative indicators of the papermilling behavior
The number of publications for each year and the number of citations in the same year are two time series, and we can investigate their properties and mutual relationship. To quantify the previous observations, the following indicators can be used when examining the publication activity of researchers:
  • correlation between the publications counts and the citations counts; high correlation coefficient (above 0.7–0.8) indicates possible papermilling;
  • the delay between the citations counts and the publications counts; if the correlation is high and the delay is small (less than 2 years), then this indicates possible papermilling;
  • large maximum number of publications in one year; unusually large number of publications in one year, with increasing trend, indicates possible papermilling;
  • low integrity index (see the next Section 6).
6 Integrity index
The main goal of a conscientious scientist is to discover new knowledge, and not to increase the number of his/her publications. In this respect, it is worth recalling Theodore von Kármán’s approach to comparing scientists [6, page 4]:
“If you define a great scientist as a man with great ideas, then you will have to rate Einstein first – he had four great ideas. In the history of science perhaps only Sir Isaac Newton is ahead of Einstein, because he had five or six ideas.
All the other major scientists of our age are associated with just one, or at the most two great ideas. In my case I have had three great ideas. Maybe more. Yes, perhaps three and a half great ideas.”
The growing numbers of published papers make increasingly difficult to distinguish the valuable, novel, and original material from material that just increases the number of published papers.
The motivation for introducing the h-index by J. E. Hirsch [7] was to focus on the publications which received sufficient attention of the related scientific community reflected in citations, and not to focus on large numbers of publications that did not draw notable attention. However, the papermills allow a fast increase of the h-index by generating a large number of worthless papers and citations to such papers, also in worthless articles; this is the aforementioned snowball effect.
The solution can be introducing a quantitative indicator, which would implement von Kármán’s idea by lowering the motivation for increasing the number of publications by penalizing the unnecessary increase of the number of publications.
As such indicator, I suggest the ratio of the h-index and the total number of publications. This can be considered as a kind of a signal-to-noise ratio, and can be called the integrity index and denoted as “$I$-index”.
If one author with h-index $H=32$ has 64 papers indexed in the Web of Science and the other author with the same h-index $H=32$ has 320 papers, then the integrity index of the first author is $I_{1}=32/64=0.5$ and the integrity index of the second author is $I_{1}=32/320=0.1$. This means the higher probability that the first author focuses on publishing novel original results and not on just publishing another paper, and the lower probability of papermilling compared to the second author.
In the example discussed in Section 4, the integrity index of researcher R1 is $I_{1}=86/196=0.44$, while the integrity index of researcher R2 is $I_{2}=34/384=0.08$. The integrity index can be used to differentiate conscientious scientists from the possible papermillers.
The values of the $I$-index of conscientious scientists are typically higher than 0.2, and in the case of papermilling the integrity index is very low compared to scientists with the similar values of h-index.
7 The MATLAB function
A function for MATLAB, “researcherprofilewos.m”, [8] has been created to simplify analysis of the data obtained from the Web of Science. The number of publications for each year and the number of citations for each year are just two times series.
This function takes an Excel file obtained from the Web of Science, and produces the following basic outputs:
  • coefficient of linear correlation between publications and citations;
  • delay between citations and publications (only meaningful if the correlation is strong);
  • maximal number of publications in one year;
  • h-index (taken from the WoS);
  • $I$-index (integrity index);
  • total number of publications;
  • total citations count.
and the following additional information:
  • starting year for publications;
  • minimal number of publications in one year;
  • average number of publications in one year;
  • average number of citations per publication.
The key output is the coefficient of linear correlation between publications and citations. Strong correlation means that the citation dynamics is strongly linked to the publication dynamics. In the case of researcher R2 the coefficient of linear correlation between publications and citations is $r_{2}=0.94$, and there is no delay between the growth of the number of publications and the number of citations. For researcher R1, the correlation coefficient is small negative number $r_{1}=-0.23$, which means that R1’s citations counts grow despite decreasing publications counts.
8 Papermilling behavior for promotions?
There are various reasons for papermilling, and being listed in the lists of highly-cited researchers is only one of them. The other reason is financial, because many universities link salaries and other financial incentives to the number of publications and citations. One of the most widespread reasons is pursuing possible promotion. Usually a candidate aspiring for promotion has to have a certain number of publications and citations in journals and conferences indexed in Web of Science or Scopus databases. The desire to achieve a promotion in a short time encourages some researchers to engage in papermilling, because doing conscientious research and waiting for the eventual future citation impact delays the desired promotion.
Two probable examples of such kind are shown in Fig. 3 (mathematics, Czech Republic), and Fig. 4 (mixed bag of environmental science, business, etc., Slovakia). with high correlation between publications and citations.
The damage done by the promotion-motivated papermilling should not be underestimated. This type of behavior is contagious and may strongly affect colleagues and students, pushing them to similar wrong and unethical practices. In addition, the promotion-motivated papermilling frequently helps such papermillers to get nto influential positions or committees where they make important decisions about hiring, grants, and promotions of honest and conscientious researchers – and also of researchers similar to themselves.
Figure 3: Papermilling for promotion: example 1 (mathematics, CZ).
Figure 4: Papermilling for promotion: example 2 (environmental science, SK).
9 Papermilling behavior outside mathematics
The exclusion of the field of mathematics from the Clarivate’s lists of highly cited researchers does not mean that other fields are spared of papermilling, as the following two examples from Slovakia illustrate.
In Fig. 5 a well-documented example from the field of plant science is shown, with very low performance in 1989–2013 and rapid growth after that: 11 papers in 2017, 14 in 2018, 19 in 2019, 33 in 2020, peaking at 90 papers in 2021, accompanied by the growth of citations counts from roughly 30 citations in 2013 to roughly 3100 citations in 2022. The number of co-authors is 1251, from 56 countries.
Figure 5: Papermilling in plant science.
Figure 6: Papermilling in engineering.
In Fig. 6 an example from the field of engineering is shown, with very low performance until 2016, and extremely rapid growth after that: 3 papers and 23 citations in 2017, 11 papers and 56 citations in 2018, 32 papers and 307 citations in 2019, more than 13 papers and around 1300 citations in 2020, more than 110 papers and more than 2500 citations in 2021, 96 papers and more than 4300 citations in 2022. The number of co-authors is 1084, coming from 63 countries.
10 Conclusion
It is obvious that the evaluation of the quality of research performance cannot be based on a single number; it does not matter if that number is the h-index, or the number of publications, or the number of citations, or the integrity index proposed in this paper, or any other single quantitative indicator.
However, there are some standard patterns typical for conscientious researchers, focused on the advancement of knowledge, and for researchers exhibiting the papermilling behavior, focusing on the numbers of publications and citations which give them advantages and benefits.
The main conclusion is that the papermilling behavior is reflected in the strong correlation between publications counts and citations counts, without standard delay between publications and citations time series; in the persistent growth of the number of publications in one year, leading to abnormally large number of articles in one year; in the low integrity index, proposed in this article; in the increasing number of co-authors along with the increase of the number of publications and citations; and with increasing number of countries where the co-authors reside. The simultaneous presence of several of these indicators means the high probability that a researcher is involved in some kind of papermilling, although there might be singular, or exceptional, cases, where the final conclusion after additional thorough examination can be different.
The detailed analysis of the dynamics of the changes of the number of co-authors related to papermilling behavior requires the change of the format of the data exported from the Web of Science database (or eventually from the Scopus database, which also can be used for such studies).
Acknowledgments
The author is supported by grants APVV-22-0508, VEGA 1/0674/23, KEGA 006TUKE-4/2024, and ARO W911NF-22-1-0264.